OpenStudy (anonymous):

The sum of two consecutive even integers is 226. What is the larger integer?

OpenStudy (amistre64):

226 :)

OpenStudy (amistre64):

n + n+2 = 226 2n + 2 = 226 2n = 224 n = 112; so n+2 is larger

OpenStudy (amistre64):

another way to do it is simply divide it in half to see what the middle of the road is; then add 1 and subtract 1

OpenStudy (anonymous):

so it's 114? :)

OpenStudy (amistre64):

im gonna say that it is, yes :)

OpenStudy (anonymous):

thannkkkyou :)

OpenStudy (anonymous):

(n-2)+n=226 n=114